OpenStudy (anonymous):

Find the exact solution of the following quadratic equation by using the Quadratic Formula. 9x^2-6x-4=-5

OpenStudy (anonymous):

\[9 x^2-6 x-4+5=0 \]\[9 x^2-6 x+1=0\]The LHS can be factored.\[(3 x-1)^2=0 \]\[\left\{x=\frac{1}{3},x=\frac{1}{3}\right\} \]